OpenStudy (anonymous):

Which of the following ordered pairs is the solution to the system of linear equations? x plus y equals fourteen, Two x plus y equals twenty-four

OpenStudy (anonymous):

Use the substitution method. y = 3x - 1 into x + 2y = 12 = x + 2(3x - 1) = 12 x + 6x - 2 = 12 (distribute 2 into 3x - 1) 7x - 2 = 12 (combine like terms) 7x = 14 (add 2 to both sides) x = 2 (divide by 7) Now that you have determined x -- y = 3x - 1 y = 3(2) - 1 y = 6 - 1 y = 5 So, your ordered pair is (2, 5) answer from yahoo answers

OpenStudy (anonymous):

oh my gosh thank you

OpenStudy (anonymous):

but i just looked at the multiple choice and their is no ordered pairs with those numbers.

OpenStudy (anonymous):

your welcome my pleasure

OpenStudy (anonymous):

ok just ignore the pair